# A Josephson Parametric Oscillator-Based Ising Machine

### 10 Dec 2023 v2

@@ 1 Introduction
The proposed structure offers advantages in terms of low power consumption and scalability, enabling the integration of single flux quantum (SFQ) based control and readout circuits on the chip. Similar circuits were proposed in [23] for qubit control.
@@ 2 Ising Machine
The physical bits form a 2-D triangular configuration, with $N-1$ bits in the base and decreasing levels above, ending in a single bit at the apex. An additional row of $N-2$ fixed physical qubits with a spin value of ’1’ completes the architecture. The connections between the physical bits are limited to nearest neighbors, resulting in $(N-1)(N-2)/2$ closed paths of length 4, known as tiles. Each tile represents a four-spin interaction pattern. Fig. 1 illustrates an example tile structure employed within the LHZ architecture with four logical JPOs, and two ancillary JPOs enforcing the constraints $C_{l}$. The constant term is added to ensure term $C_{l}$ is always negative.
Figure 1: A general structure of the JPO-based tile. $S_{1}$ to $S_{4}$ (the equivalent of n.e.s.w. in the LHZ) are the JPO-based spin nodes that are coupled together, and $S_{a1}$ and $S_{a2}$ are JPO-based ancilla spins which are used to impose the constraint $C_{l}$. $C_{cnst}$ is the offset value added to ensure that $C_{l}$ is always negative in the tile structure.
@@ 3.1 Josephson Parametric Oscillator (JPO)
where $L_{JJ}$ is the inductance observed across the JJ, and $\Phi_{0}$ is the quantum of magnetic flux. Here $I_{C}$ is the critical current that can be described as $I_{C0}\cos(\varphi)$ where $\varphi$ is the junction’s phase and $I_{C0}$ is the critical current at a zero phase value. The nonlinear relationship between phase and current in the JJ makes it an excellent candidate for nonlinear circuits and oscillators in different devices such as superconductor quantum interference devices (SQUIDs) and qubits [26, 27, 28].
@@
[…]
@@ 3.2 Tile (Plaquette)
Table 1 illustrates the stable minimum energy states achieved in the circuit that depend on the values the ancillary JPOs settle into. When the ancillary settle into the $|00\rangle$ state, they force the physical JPOs to go to the $|1111\rangle$ state. When we set the ancillas to $|01\rangle$ or $|10\rangle$, only two logical JPOs can settle into the $|1\rangle$ state, and finally, in case the ancillas are settled at the $|11\rangle$ state, they force the physical JPOs to go to the $|0000\rangle$ state. This satisfies the ground state condition of the LHZ tile, where only even numbers of similar states should exist in a single tile. Note that in this table, only the valid states based on (3) have been provided.
@@ 4.1 Analytical modeling of Ising Hamiltonian
where $J_{i}$ represent external field parameters, while $C_{t}$ represents the interaction arising from the coupler ($C_{t}=\sigma^{x}_{5}\times J_{a}+\sigma^{x}_{6}\times J_{a}+J_{,C}$), and the terms $\sigma^{(z)}$ correspond to Pauli’s z matrix. Additionally, we incorporate $U_{N}$ as a noise component generated by multiplying a thermal coefficient by a pseudo-random energy distribution. The developed Python code for the tile’s Ising Hamiltonian calculation is openly available on GitHub [^1] Python code available at github.com/SasanRazm/IsingQutip.
The minimum energy states of the tile are investigated by introducing all the possible states and measuring the minimum energy value of the tile based on its Hamiltonian equation introduced in equation 9. For simulation, the external field was normalized to 1, and the interaction between the spins was $\pm$0.2 with all the positive and negative combinations. Then, based on the number of times the circuit is settled on a state, we calculate the distribution probability. The result for the probability of the states is shown in Fig. 5. Here, the tile settles on the states that the number of 1’s are even, as shown in Table 1. The difference between the probabilities of the states is due to the noise we have considered in the simulations.
@@ 4.2 Analog simulation setup
With the proposed design, higher frequency values can be achieved. While higher frequencies result in a smaller tile layout and better noise performance, we chose this value to align with our plan of utilizing on-chip control and readout circuits. The control circuits will use SFQ pulses to generate the microwave pump signal. The frequency of the SFQ pulses should be about five times the pump frequency. After filtering to get a clear sinusoidal signal and attenuation to match the noise level and avoid quasi-particle poisoning, the signals are applied to the tiles via coupled inductances.
However, for analog simulation, we did not use control circuits, and signals were applied directly via voltage sources. We employed the DC flux by introducing it to the SQUID loop through a DC voltage source via coupled inductance with the SQUID loop inductances. To enhance the accuracy of the simulation, a resistor with thermal noise was incorporated into the voltage source. A similar approach was adopted for applying AC pump flux to the SQUID loop. The presence of coupled inductances can impact the behavior of the SQUID loop and resonator, and as a result, their influence needs to be considered in the overall circuit design.
@@
Simulating a tile is more complicated than JPO since the inputs should be applied to 6 different JPOs. For a tile, the pump voltage phases were applied based on the $J_{i}$ values. Then the DC flux was increased to destabilize the circuit, gradually decreasing to the design point. At this point, the SQUID voltage of each JPO is measured, and the phase of it was either at 0 or $\pi$. This process was done a thousand times to achieve the statistical distribution of the states.

### 12 Dec 2023 v3

@@ 2 Ising Machine
The physical bits form a 2-D triangular configuration, with $N-1$ bits in the base and decreasing levels above, ending in a single bit at the apex. An additional row of $N-2$ fixed physical qubits with a spin value of ’1’ completes the architecture. Each tile represents a four-spin interaction pattern. Fig. 1 illustrates an example tile structure employed within the LHZ architecture with four logical JPOs, and two ancillary JPOs enforcing the constraints $C_{l}$. The constant term is added to ensure term $C_{l}$ is always negative.
Figure 1: A general structure of the JPO-based tile. $S_{1}$ to $S_{4}$ are the JPO-based spin nodes that are coupled together, and $S_{a1}$ and $S_{a2}$ are JPO-based ancilla spins which are used to impose the constraint $C_{l}$. $C_{cnst}$ is the offset value added to ensure that $C_{l}$ is always negative in the tile structure.
@@ 3.1 Josephson Parametric Oscillator (JPO)
where $L_{JJ}$ is the inductance observed across the JJ, and $\Phi_{0}$ is the quantum of magnetic flux. The nonlinear relationship between phase and current in the JJ makes it an excellent candidate for nonlinear circuits and oscillators in different devices such as superconductor quantum interference devices (SQUIDs) and qubits [26, 27, 28].
@@
Figure 3: Resonator frequency calculation result with different applied DC currents. The material is Nb, and the characteristics are chosen to match the MIT LL SFQee5 process.
@@ 3.2 Tile (Plaquette)
Table 1 illustrates the stable minimum energy states achieved in the circuit that depend on the values the ancillary JPOs settle into. When the ancillary settle into the $|00\rangle$ state, they force the physical JPOs to go to the $|1111\rangle$ state. When we set the ancillas to $|01\rangle$ or $|10\rangle$, only two logical JPOs can settle into the $|1\rangle$ state, and finally, in case the ancillas are settled at the $|11\rangle$ state, they force the physical JPOs to go to the $|0000\rangle$ state. This satisfies the ground state condition of the LHZ tile, where only even numbers of similar states should exist in a single tile.
@@ 4.1 Analytical modeling of Ising Hamiltonian
where $J_{i}$ represent external field parameters, while $C_{t}$ represents the interaction arising from the coupler ($C_{t}=\sigma^{x}_{5}\times J_{a}+\sigma^{x}_{6}\times J_{a}+J_{,C}$), and the terms $\sigma^{(z)}$ correspond to Pauli’s z matrix. Additionally, we incorporate $U_{N}$ as a noise component generated by multiplying a thermal coefficient by a pseudo-random energy distribution.
The minimum energy states of the tile are investigated by introducing all the possible states and measuring the minimum energy value of the tile based on its Hamiltonian equation introduced in equation 9. Then, based on the number of times the circuit is settled on a state, we calculate the distribution probability. The result for the probability of the states is shown in Fig. 5. Here, the tile settles on the states that the number of 1’s are even, as shown in Table 1.